A simple random sample of size

n=49

is obtained from a population with

μ=73

and

σ=21.

​(a) Describe the sampling distribution of

x.

​(b) What is

P x>78.7​?

​(c) What is

P x≤66.25​?

​(d) What is

P 70.6<x<80.2​?
Answer #1

Giron na ug M=73 221 by Centre Limit Theorem (CLT) sampling distribution of it is normally distributed with mean (4x) and stab] P ( x > 78.9) P I-lx > 78.7-73 plzs lig - I-plz<1.9) = -0.9013 = 10.0287 Answer i-0.0289 13) e] PC & <66.25) = P(Z< 66.25-3 a] P(70.6<$<80.2) = P(10.6=73 <75 80,2,-73 = P(-0.84252.4) p(z<2.4) -P(240.8) 8(2624)*P(2¢0.87] =P(2.82.4) -AP (2508) 0.991